# Self-Admitted Technical Debt in LLM Software: An Empirical Comparison with ML and Non-ML Software

### v2

@@ IV-F SATD Prevalence Metrics
File-Level Prevalence (Diffusion). This metric measures the spread of technical debt across the codebase, indicating how widely SATD affects the project’s architecture, as follows:
@@ IV-G Statistical Analysis Methods
Wilcoxon Rank-Sum Test: We use the Wilcoxon rank-sum test (Mann-Whitney U test) [36] to compare SATD prevalence between repository types. This non-parametric test is suitable for our non-normally distributed as SATD percentages with varying sample sizes and is robust to outliers [23].
Effect Size (Cohen’s d): While p-values indicate statistical significance, Cohen’s $d$ [5] quantifies the practical magnitude of differences. We interpret effect sizes following Cohen’s conventions: $d<0.2$ (negligible), $0.2\leq d<0.5$ (small), $0.5\leq d<0.8$ (medium), $d\geq 0.8$ (large). This distinction is crucial as large sample sizes can yield statistical significance even for trivially small differences [35].
Bonferroni Correction: To control the family-wise error rate in multiple pairwise comparisons (LLM vs. ML, LLM vs. Non-ML, and ML vs. Non-ML), we apply Bonferroni correction [7], lowering the significance threshold $\alpha$ from 0.05 to 0.017 for three comparisons or three comparisons. We prioritize controlling false positives to ensure that reported differences between repository types reflect genuine patterns rather than statistical artifacts from multiple testing.”
Survival Analysis: For temporal analyses (RQ4), we employ Kaplan-Meier estimation [17] to model SATD introduction and removal times, with log-rank tests [24] for comparing survival curves across repository types. This approach properly handles censored data (SATD still present at study endpoint).
@@ V-A RQ1: SATD Prevalence Across Repository Types
Table I presents SATD prevalence results measured as the ratio of SATD comments to total comments. We found 3.95% SATD in LLM repositories, 4.10% in ML repositories, and 3.23% in non-ML repositories.
@@
The prevalence difference between ML and non-ML repositories is $1.27x$, compared to $2x$ as reported by Bhatia et al. [4]. Wilcoxon rank-sum tests confirm significant differences between all repository pairs (all $p<0.005$), exceeding the Bonferroni-corrected significance threshold of $\alpha=0.017$ for three comparisons. Effect sizes indicate moderate practical significance for ML vs. non-ML (Cohen’s d = 0.42) and LLM vs. non-ML (d = 0.38) comparisons, but minimal difference between ML-LLM repositories (d = 0.04). The comparable SATD prevalence between LLM (3.95%) and ML (4.10%) repositories challenges our initial hypothesis that LLM systems would accumulate substantially more technical debt due to their architectural complexity, external API dependencies, and fast-paced ecosystems. This unexpected similarity (d = 0.04) suggests that LLM development has benefited from lessons learned in earlier ML systems, adopting disciplined practices from inception rather than repeating the experimental chaos that characterized early ML projects.
However, this similarity must be interpreted considering the different temporal contexts of our repository cohorts. The ML repositories, sampled from Bhatia et al.’s 2015-2021 dataset, have had several years to mature and potentially undergo cycles of debt accumulation and removal by the time of our analysis. In contrast, LLM repositories are relatively young (post-2022), capturing them in their initial development phase. The post-2022 LLM ecosystem has been characterized by a high-pressure, rapid-development, with teams racing to achieve state-of-the-art results and capture market share. This environment incentivizes shortcuts and quick iterations that would bias the LLM cohort toward accumulating more technical debt in shorter time frames.
Practical Implication. The statistically similar levels of technical debt between LLM and ML projects (d = 0.04) suggest two opposing forces: the stabilizing effect of ML project maturity versus the rapid “gold rush” development of LLMs. The comparable SATD levels may indicate either that LLM projects accumulate debt more quickly, reaching ML-level prevalence within 2–3 years, or that modern LLM development benefits from established ML practices. The narrowing gap between ML and non-ML repositories ($1.27x$ vs. $2x$ in 2021 [4]) reflects broader AI development maturation, likely aided by standardized frameworks like PyTorch Lightning and Hugging Face Transformers.
@@ V-B RQ2: Types of SATD in LLM-based Systems
We analyzed 377 sampled SATD instances from LLM repositories to identify debt categories and extend existing taxonomies. Figure 1 presents the distribution of SATD types, revealing that while traditional categories remain dominant, 18% of classified SATD represents novel LLM-specific patterns.
@@ V-B1 Extending Existing Taxonomies for Modern AI Systems
Our analysis revealed key limitations in applying traditional SATD taxonomies to LLM systems. While Bhatia et al. [4] successfully extended Bavota and Russo’s taxonomy [3] with two ML-specific categories (Configuration Debt, Inadequate Testing), even these extensions prove insufficient for LLM systems. The original taxonomy assumes object-oriented architectures with violations like poor encapsulation, which is suitable for Bhatia et al.’s OO-heavy ML repositories but incompatible with modern LLM systems that use procedural logic, distributed runtimes, and hardware-specific pipelines [32]. This architectural shift necessitated substantial taxonomy expansion: while Bhatia et al. needed only minor additions for ML debt, 18% of LLM SATD required entirely new categories. Their Configuration Debt captured hyperparameter tuning but not external API dependencies (Model-Stack Workaround Debt); their taxonomy addressed static ML pipelines but not dynamic ecosystem changes (Model Dependency Debt). This evolution from minor extensions to major new categories reflects the fundamental architectural leap from traditional ML to LLM development.
As a result, many SATD instances labeled as design issues are not OO violations but architectural problems, such as misplaced responsibilities (TODO(matt): move into LLMConfig) or inconsistent data handling (TODO: Clean this up to only use one type). To account for this, we broadened Design Debt to include any architectural, procedural, or structural deficiencies that harm maintainability. This expanded definition captured 14.6% of classified SATD, making it the most common category in our sample.
We also found that traditional Workaround Debt assumes a unified software stack under developer control. LLM systems violate this assumption because they depend on external model stacks that often lack features or contain bugs [40]. Comments such as “HACK: No official tokenizer available for Claude 3” and “This is a workaround for the TPU SPMD mode” illustrate compensations for external infrastructure deficiencies rather than internal code issues.
@@ V-B2 New Debt Categories in the LLM Ecosystem
Beyond taxonomy extensions, we identified three categories of SATD specific to LLM development, representing 18% of classified instances. Model-Stack Workaround Debt (7.4%) represents hacks and compatibility fixes for external ML infrastructure limitations. This category, distinct from traditional workarounds, reflects the reality that LLM developers must constantly adapt to bugs and missing features in fast-paced model APIs, tokenizers, and runtime environments they cannot control. Model Dependency Debt (3.7%) captures deferred updates waiting for upstream library evolution. Comments such as “TODO: Remove once accelerate is updated” and “TODO: will remove after torch xpu 2.9 supports uuid” reveal a dependency-driven debt pattern where code remains provisional until external frameworks provide required features. Unlike traditional dependency management, resolution depends entirely on external release cycles. Performance Optimization Debt (4.0%) represents deferred hardware-specific optimizations requiring specialized expertise. Comments reference GPU kernel engineering (TODO: optimize with triton kernels), memory layout adjustments, and FlashAttention integration (an attention mechanism optimization for transformers). This differs from traditional performance debt, which involves algorithmic improvements, as it requires deep knowledge of hardware acceleration and kernel programming that is often absent from application teams.
The emergence of LLM-specific debt categories shows how ecosystem volatility shapes technical debt in modern AI systems. The dominance of Model-Stack Workaround Debt, the largest LLM-specific category, indicates that external infrastructure instability drives much of this debt. Developers expend significant effort compensating for limitations in APIs and frameworks beyond their control, a pattern less common in traditional software where most dependencies are stable. Despite these novel challenges, traditional categories remain prevalent. Requirements Debt (11.9%) and Code Debt (10.6%) persist alongside LLM-specific issues, showing that core software engineering challenges transcend paradigm shifts. The high proportion of “Undefined” SATD (comments containing only “TODO”) suggests that debt documentation practices lag behind system complexity.
@@
Answer to RQ2: LLM repositories exhibit 18% novel SATD types: Model-Stack Workaround (7.4%), Model Dependency (3.7%), and Performance Optimization Debt (4.0%). While traditional debt categories like Design (14.6%) and Requirements (11.9%) remain dominant, these LLM-specific types reflect unique challenges in external infrastructure dependencies and hardware optimization not present in traditional or pre-LLM ML systems.
@@ V-C RQ3: Distribution of SATD Across LLM Development Pipeline Stages
We mapped SATD instances to the LLM development pipeline stages defined by Hu et al. [14] to understand where technical debt concentrates in modern LLM workflows. Figure 2 presents the distribution across five pipeline stages plus a General category for cross-cutting concerns.
@@
The concentration of SATD in Deployment and Monitoring (30% of stage-specific debt) highlights the operational difficulty of serving LLMs at scale. This stage requires sophisticated optimizations, such as quantization, kernel tuning, memory scheduling, and parallel serving, thus creating many opportunities for provisional solutions. Some comments acknowledge deferred optimizations (TODO: implement dynamic batching), temporary workarounds for serving constraints (HACK: fixed batch size until memory profiling complete), and configuration shortcuts taken under production pressure.
This distribution contradicts assumptions that model logic drives technical debt; instead, operational infrastructure emerges as the primary debt source. The complexity of operationalizing LLMs, balancing latency, throughput, and resource utilization, forces teams to accept provisional solutions that often become permanent, with many SATDs referencing integration challenges with serving frameworks and deferred optimizations awaiting specialized expertise.
Pretraining (23%) represents the second major concentration, with SATD referencing missing checkpointing logic, incomplete parallelization, and deferred memory optimizations. The resource-intensive nature of pretraining creates pressure to maintain forward progress rather than address technical debt.
Alignment (4%) and Evaluation (8%) show minimal SATD, consistent with their short-running jobs and modest resource requirements. Their debt involves missing features (TODO: add ROUGE metric) rather than architectural issues. This pattern suggests that modularity and operational simplicity act as natural barriers to debt accumulation, as evaluation and alignment scripts can be modified independently, making debt both less likely to accumulate and easier to address.
@@
V-D RQ4: Temporal SATD Characteristics of Across Repository Types
We analyzed the complete git history of all repositories to understand when SATD appears and how long it persists across different software paradigms. Our survival analysis methodology, detailed in the technical box below, tracks both introduction timing and persistence patterns. Table III summarizes the temporal characteristics from our analysis of over 5 million comment events.
@@ V-D RQ4: Temporal SATD Characteristics of Across Repository Types
TABLE III: SATD temporal characteristics (* refers to total additions and removals across all commits)
@@ V-D1 LLM Projects: Delayed Introduction, Persistent Accumulation
LLM repositories exhibit a two-phase development pattern. They remain SATD-free for a median of 492 days, far longer than ML projects (204 days) but shorter than traditional software (1,005 days), reflecting a prolonged infrastructure-building period before complexity forces shortcuts. Once SATD appears, it tends to persist. With a 553-day median lifespan and the lowest removal rate (49.1%), debt becomes structurally embedded. Tight coupling across prompts, APIs, and orchestration layers creates architectural lock-in, making refactoring costly, unlike ML systems where experimental components can be replaced with far less risk. File-level analysis (Table IV) reinforces this pattern: LLM files wait 1,144 days (median) before first SATD appears, suggesting careful initial development followed by inevitable compromise as complexity grows.
@@ V-D2 Repository Maturation: Understanding the $44x$ Shift
We observed a major temporal shift in ML repositories: the median time to first SATD increased from 10 days (Bhatia et al., 2021) to 441 days in our 2025 analysis, a $44x$ change. This reflects repository maturation rather than methodological differences. Early ML repositories (around the year 2021) were in experimental phases with rapid file creation and immediate SATD as developers prototyped [27]. By 2025, many of these experimental files have been removed or stabilized. The remaining files form a mature core that accumulated SATD only after multiple development cycles. This demonstrates that SATD behavior evolves over the project lifecycle. The result highlights a methodological challenge: cross-sectional studies capture only a snapshot and can misrepresent long-term debt patterns. SATD analysis requires a temporal perspective, as the same repository can exhibit fundamentally different behaviors at different stages of maturity.
@@ V-E RQ5: Predictors of Long-Lasting SATD
Following Bhatia et al. [4], we trained Random Forest classifiers to identify what likely makes some technical debts last longer than others. We analyzed commit-level features (lines added, lines modified, files changed) and file-level features (lines of code, complexity metrics, token counts) to predict debt persistence. To find predictive patterns, the models classify each debt as (1) long-lasting debt (surviving longest 25%), or (2) quickly-removed debt (removed fastest 25%). Table VI presents the model performance and top predictors for each repository type.
@@
Our models predict long-lasting debt with 96% accuracy in ML repositories, 82% in LLM repositories, and 74% in traditional software. ML projects follow similar experimental workflows, making debt patterns consistent and predictable. LLM projects mix experimentation with production engineering, creating more varied debt patterns. Traditional software has the most diverse debt sources—user interfaces, databases, business logic—making prediction hardest. By examining feature importance (Table VI), we see that the predictors reveal why debt persists across different systems. In ML repositories, large commits (hundreds of lines added or modified) create the most persistent debt. When researchers run large experiments or refactor entire pipelines, the resulting debt often becomes permanent. LLM repositories show a different pattern: debt in large files (by lines of code or token count) persists, reflecting that large inference pipelines and serving infrastructure are difficult to refactor once deployed. Traditional software relies on complexity metrics. When code becomes complex (high cyclomatic complexity) or files grow large, debt becomes harder to understand and fix, thus persists. One universal pattern emerges: large commits create lasting debt regardless of project type. When teams rush to ship features or fix urgent problems, they create debt that rarely gets cleaned up. Teams can apply these findings immediately.Since commit size strongly predicts persistent debt, automated tools could flag large commits (>500>500 lines) for extra review. For LLM projects, review should focus on large infrastructure files where debt most likely becomes permanent. Varying prediction accuracy guides tool development. ML teams can use simple commit-size rules with high confidence. LLM and traditional software teams need sophisticated multi-factor approaches. SATD tools must adapt to each domain rather than rely on universal approaches.

### v3

@@ IV-F SATD Prevalence Metrics
File-Level Prevalence (Diffusion). This metric measures the spread of technical debt across the codebase, indicating how widely SATD affects the project’s architecture as follows:
@@ IV-G Statistical Analysis Methods
Wilcoxon Rank-Sum Test: We use the Wilcoxon rank-sum test (Mann-Whitney U test) [35] to compare SATD prevalence between repository types. This non-parametric test is suitable for our non-normally distributed SATD percentages with varying sample sizes and is robust to outliers [21].
Effect Size (Cohen’s d): While p-values indicate statistical significance, Cohen’s $d$ [4] quantifies the practical magnitude of differences. We interpret effect sizes following Cohen’s conventions: $d<0.2$ (negligible), $0.2\leq d<0.5$ (small), $0.5\leq d<0.8$ (medium), $d\geq 0.8$ (large). This distinction is crucial, as large sample sizes can yield statistical significance even for trivially small differences [34].
Bonferroni Correction: To control the family-wise error rate in multiple pairwise comparisons (LLM vs. ML, LLM vs. Non-ML, and ML vs. Non-ML), we apply Bonferroni correction [6], lowering the significance threshold $\alpha$ from 0.05 to 0.017 for three comparisons. We prioritize controlling false positives to ensure that reported differences between repository types reflect genuine patterns rather than statistical artifacts from multiple testing.”
Survival Analysis: For temporal analyses (RQ4), we employ Kaplan-Meier estimation [15] to model SATD introduction and removal times, with log-rank tests [22] for comparing survival curves across repository types. This approach properly handles censored data, which occurs when SATD remains unresolved at the study endpoint and we observe only a minimum survival time rather than the actual removal time.
@@ V-A RQ1: SATD Prevalence Across Repository Types
Table I presents SATD prevalence results (density) measured as the ratio of SATD comments to total comments. We found 3.95% SATD in LLM repositories, 4.10% in ML repositories, and 3.23% in non-ML repositories.
@@
The prevalence difference between ML and non-ML repositories is $1.27x$, compared to $2x$ as reported by Bhatia et al. [3]. Wilcoxon rank-sum tests confirm significant differences between all repository pairs (all $p<0.005$), exceeding the Bonferroni-corrected significance threshold of $\alpha=0.017$ for three comparisons. Effect sizes indicate moderate practical significance for ML vs. non-ML (Cohen’s d = 0.42) and LLM vs. non-ML (d = 0.38) comparisons, but minimal difference between ML-LLM repositories (d = 0.04). The comparable SATD prevalence between LLM (3.95%) and ML (4.10%) repositories challenges our initial hypothesis that LLM systems would accumulate substantially more technical debt due to their architectural complexity, external API dependencies, and fast-paced ecosystems. This similarity (d = 0.04) suggests that LLM development has benefited from lessons learned in earlier ML systems, adopting disciplined practices from inception rather than repeating the experimental chaos that characterized early ML projects.
However, this similarity must be interpreted considering the different temporal contexts of our repository cohorts. The ML repositories, sampled from Bhatia et al.’s 2015-2021 dataset, have had several years to mature and potentially undergo cycles of debt accumulation and removal by the time of our analysis. In contrast, LLM repositories are relatively young (post-2022), capturing them in their initial development phase. The post-2022 LLM ecosystem has been characterized by a high-pressure, fast-moving development, with teams racing to achieve state-of-the-art results and capture market share. This environment incentivizes shortcuts and quick iterations that would bias the LLM cohort toward accumulating more technical debt in shorter time frames.
Practical Implication. The statistically similar levels of technical debt in LLM and ML projects (d = 0.04) reflect two opposing forces: the stabilizing effect of mature ML practices and the rapid “gold rush” pace of LLM development. The comparable SATD levels may indicate LLM projects accumulate debt faster, matching ML prevalence within 2–3 years, or that modern LLM development benefits from established ML practices. The narrowing gap between ML and non-ML repositories ($1.27x$ vs. $2x$ in 2021 [3]) likely signals broader AI maturation, supported by standardized frameworks such as PyTorch Lightning and Hugging Face Transformers.
@@ V-B RQ2: Types of SATD in LLM-based Systems
We analyze 377 SATD instances from our LLM repositories to identify debt categories and extend existing taxonomies. Figure 1 presents the distribution of SATD types, revealing that although traditional categories remain dominant, 15% of classified SATD comprises novel LLM-specific patterns.
@@ V-B1 Extending Existing Taxonomies for Modern AI Systems
Our analysis shows that existing SATD taxonomies do not adequately capture debt in LLM systems. Bhatia et al. [3] extended Bavota and Russo’s taxonomy [2] with two ML-specific categories (Configuration Debt, Inadequate Testing), but these remain insufficient for LLM systems. The original taxonomy assumes object-oriented architectures (e.g., poor encapsulation), which fits Bhatia et al.’s OO-heavy ML repositories but not modern LLM systems that rely on procedural logic, distributed runtimes, and hardware-specific pipelines [31]. This architectural shift required substantial taxonomy expansion: although Bhatia et al. made only minor ML-specific additions, 22.4% of LLM SATD in our study required entirely new categories. Their Configuration Debt captured hyperparameter tuning but not external API dependencies (Model-Stack Workaround Debt), and their taxonomy models static ML pipelines but not dynamic ecosystem changes (Model Dependency Debt). This move from minor extensions to many new categories reflects a fundamental architectural gap between LLM and traditional ML development.
As a result, many SATD instances labeled as design issues are not OO violations but architectural problems, such as misplaced responsibilities (“TODO(matt): move into LLMConfig”) or inconsistent data handling (“TODO: Clean this up to only use one type”). To account for this, we broaden Design Debt to cover any architectural, procedural, or structural deficiencies that harm maintainability. This expanded definition captured 14.6% of classified SATD, making it the most common category in our sample. We also found that traditional Workaround Debt assumes a unified software stack under developer control. LLM systems violate this assumption because they depend on external model stacks that often lack features or contain bugs [39]. Comments such as “HACK: No official tokenizer available for Claude 3” and “This is a workaround for the TPU SPMD mode” illustrate compensations for external infrastructure deficiencies rather than internal code issues.
@@ V-B2 New Debt Categories in the LLM Ecosystem
Beyond taxonomy extensions, we identify three categories of SATD specific to LLM development, representing 22.4% of classified instances. Model-Stack Workaround Debt (6.3%) represents hacks and compatibility fixes for external ML infrastructure limitations. This category, distinct from traditional workarounds, reflects the reality that LLM developers must constantly adapt to bugs and missing features in fast-paced model APIs, tokenizers, and runtime environments they cannot control. Model Dependency Debt (11.2%) captures deferred updates waiting for upstream library evolution. Comments such as “TODO: Remove once accelerate is updated” and “TODO: will remove after torch xpu 2.9 supports uuid” reveal a dependency-driven debt pattern where code remains provisional until external frameworks provide required features. Unlike traditional dependency management, resolution depends entirely on external release cycles. Performance Optimization Debt (4.9%) represents deferred hardware-specific optimizations requiring specialized expertise. Comments reference GPU kernel engineering (“TODO: optimize with triton kernels”), memory layout adjustments, and FlashAttention integration (an attention mechanism optimization for transformers). This differs from traditional performance debt, which involves algorithmic improvements, as it requires deep knowledge of hardware acceleration and kernel programming that is often absent from application teams.
The prevalence of LLM-specific debt categories shows how ecosystem volatility shapes technical debt in modern AI systems. The dominance of Model-Stack Workaround Debt, the largest LLM-specific category, indicates that external infrastructure instability drives much of this debt. Developers expend significant effort compensating for limitations in APIs and frameworks beyond their control, a pattern less common in traditional software where most dependencies are stable. Despite these novel challenges, traditional categories remain prevalent. Requirements Debt (11.9%) and Code Debt (10.6%) persist alongside LLM-specific issues, showing that core software engineering challenges transcend paradigm shifts. The high proportion of “Undefined” SATD (comments containing only “TODO”) suggests that debt documentation practices lag behind system complexity.
@@
Answer to RQ2: LLM repositories exhibit 22.4% novel SATD types: Model-Stack Workaround (6.3%), Model Dependency (11.2%), and Performance Optimization Debt (4.9%). While traditional debt categories like Design (11.2%) and Requirements (23.1%) remain dominant, these LLM-specific types reflect unique challenges in external infrastructure dependencies and hardware optimization not present in traditional or pre-LLM ML systems.
@@ V-C RQ3: Distribution of SATD Across LLM Development Pipeline Stages
We map SATD instances to the LLM development pipeline stages defined by Hu et al. [13] to understand where technical debt concentrates in modern LLM workflows. Figure 2 presents the distribution across five pipeline stages plus a General category for shared infrastructure not tied to any specific pipeline stage. The concentration of SATD in Deployment and Monitoring (30% of stage-specific debt) highlights the operational difficulty of serving LLMs at scale. This stage requires sophisticated optimizations, such as quantization, kernel tuning, memory scheduling, and parallel serving, thus creating many opportunities for provisional solutions. Some comments acknowledge deferred optimizations (“TODO: implement dynamic batching”), temporary workarounds for serving constraints (“HACK: fixed batch size until memory profiling complete”), and configuration shortcuts taken under production pressure.
@@
This distribution shows that operational infrastructure, rather than model logic, drives debt accumulation; instead, operational infrastructure emerges as the primary debt source. The complexity of operationalizing LLMs, balancing latency, throughput, and resource utilization, forces teams to accept provisional solutions that often become permanent, with many SATDs referencing integration challenges with serving frameworks and deferred optimizations awaiting specialized expertise.
Pretraining (23%) is the second-largest concentration of debt, with SATD referencing missing checkpointing logic, incomplete parallelization, and deferred memory optimizations. The resource-intensive nature of pretraining creates pressure to maintain forward progress rather than address technical debt.
Alignment (4%) and Evaluation (8%) show minimal SATD, consistent with their short-running jobs and modest resource requirements. Debt in these stages involves missing features (“TODO: add ROUGE metric”) rather than architectural issues. This pattern suggests that modularity and operational simplicity act as natural barriers to debt accumulation, as evaluation and alignment scripts can be modified independently, making debt both less likely to accumulate and easier to address.
@@
V-D RQ4: Temporal SATD Characteristics Across Repository Types
We analyze the complete git history of all repositories to understand when SATD appears and how long it persists across different software paradigms. Our survival analysis methodology (detailed below) tracks both introduction timing and persistence patterns. Table III summarizes the temporal characteristics from our analysis of over 5 million comment events.
@@ V-D RQ4: Temporal SATD Characteristics Across Repository Types
TABLE III: SATD temporal characteristics (* refer to total additions and removals across all commits)
@@ V-D1 LLM Projects: Delayed Introduction, Persistent Accumulation
LLM repositories exhibit a two-phase development pattern. They remain SATD-free for a median of 492 days, far longer than ML projects (204 days) but shorter than traditional software (1,005 days), reflecting a prolonged infrastructure-building phase before complexity forces shortcuts. Once SATD appears, it persists: with a 553-day median lifespan and the lowest removal rate (49.1%), debt becomes structurally embedded. Tight coupling across prompts, APIs, and orchestration layers creates architectural lock-in, making refactoring costly, unlike ML systems, where experimental components can be replaced with far less risk. File-level analysis (Table IV) confirms this: LLM files wait 1,144 days (median) before first SATD appears, suggesting careful early development followed by inevitable compromise as complexity grows.
@@ V-D2 Repository Maturation: Understanding the $44x$ Shift
Our results show a major temporal shift in ML repositories: the median time to first SATD increased from 10 days (Bhatia et al., 2021) to 441 days in our 2025 analysis, a $44x$ increase. This reflects repository maturation rather than methodological differences. Around 2021, ML repositories were in experimental phases with rapid file creation and immediate SATD as developers prototyped [25]. By 2025, many experimental files had been removed or stabilized, leaving a mature core where SATD appears only after multiple development cycles. This demonstrates that SATD behavior evolves over the project lifecycle. The result highlights a methodological challenge: cross-sectional studies capture only a snapshot and can misrepresent long-term debt patterns. SATD analysis therefore requires a temporal perspective, as the same repository can behave very differently at different stages of maturity.
@@ V-E RQ5: Predictors of Long-Lasting SATD
Following Bhatia et al. [3], we train Random Forest classifiers to identify what likely makes some technical debts last longer than others. We analyze commit-level features (lines added, lines modified, files changed) and file-level features (lines of code, complexity metrics, token counts) to predict debt persistence. To find predictive patterns, the models classify each debt as (1) long-lasting debt (surviving longest 25%), or (2) quickly-removed debt (removed fastest 25%). Table VI presents the model performance and top predictors for each repository type.
@@
Our models predict long-lasting debt with 96% accuracy in ML repositories, 82% in LLM repositories, and 74% in traditional software. ML projects follow similar experimental workflows, making debt patterns consistent and predictable. LLM projects mix experimentation with production engineering, creating more varied debt patterns. Traditional software has the most diverse debt sources—user interfaces, databases, business logic—making prediction hardest. By examining feature importance (Table VI), we observe that the predictors reveal why debt persists across different systems. In ML repositories, large commits (hundreds of lines added or modified) create the most persistent debt. When researchers run large experiments or refactor entire pipelines, the resulting debt often becomes permanent. LLM repositories show a different pattern: debt in large files (by lines of code or token count) persists, reflecting that large inference pipelines and serving infrastructure are difficult to refactor once deployed. Traditional software relies on complexity metrics. When code becomes complex (high cyclomatic complexity) or files grow large, debt becomes harder to understand and fix, thus persists.
A universal pattern emerges: large commits create lasting debt regardless of project type. When teams rush to ship features or fix urgent problems, they create debt that rarely gets cleaned up. Teams can act on this immediately: because commit size strongly predicts persistent debt, tools should flag large commits (>500>500 lines) for extra review. In LLM projects, review should focus on large infrastructure files where debt most likely becomes permanent. Varying prediction accuracy guides tool development. ML teams can use simple commit-size rules with high confidence, while LLM and traditional software teams need sophisticated multi-factor approaches. SATD tools must adapt to each domain rather than rely on universal approaches.